# Pointer Networks Trained Better via Evolutionary Algorithms

### 10 Dec 2023 v3

@@ Abstract
Pointer Network (PtrNet) is a specific neural network for solving Combinatorial Optimization Problems (COPs). While PtrNets offer real-time feed-forward inference for complex COPs instances, its quality of the results tends to be less satisfactory. One possible reason is that such issue suffers from the lack of global search ability of the gradient descent, which is frequently employed in traditional PtrNet training methods including both supervised learning and reinforcement learning. To improve the performance of PtrNet, this paper delves deeply into the advantages of training PtrNet with Evolutionary Algorithms (EAs), which have been widely acknowledged for not easily getting trapped by local optima. Extensive empirical studies based on the Travelling Salesman Problem (TSP) have been conducted. Results demonstrate that PtrNet trained with EA can consistently perform much better inference results than eight state-of-the-art methods on various problem scales. Compared with gradient descent based PtrNet training methods, EA achieves up to 30.21% improvement in quality of the solution with the same computational time. With this advantage, this paper is able to at the first time report the results of solving 1000-dimensional TSPs by training a PtrNet on the same dimensionality, which strongly suggests that scaling up the training instances is in need to improve the performance of PtrNet on solving higher-dimensional COPs.
@@ 1 Introduction
On the other hand, one major drawback of PtrNet is the lower quality of the results compared with traditional methods, especially on larger-scale COPs, as systematically evaluated by [13] . For example, the quality gap of result between PtrNet and the well-established solver Lin-Kernighan heuristic (LKH) [14] degenerate significantly from 19.81% to 41.44% on TSP with 250 nodes and 1000 nodes [15], respectively. Although recent works have proposed more advanced models like Attention [16] and POMO [17] , they still cannot resolve the above issue. On this basis, this paper focuses on the training method of the neural networks (NNs) for COPs, rather than the model design. We initially opted for the PtrNet over the more complex models for simplicity and generality. The straightforward architecture of PtrNet provides a clear foundation for implementing and understanding the effects of EAs in training NNs, and establishes a baseline before exploring more intricate models.
@@
In order to verify the aforementioned advantages of EAs, we have designed sophisticated computational experiments on training PtrNet for TSPs. The experiments specially focus on an important training challenge. That is, existing training methods are hardly to train PtrNet effectively on 1000-dimensional instances, which are more realistic to real-world problem scales [29]. Albeit PtrNet is able to handle varied dimensional instances on its recurrent network structure, the quality of solutions deteriorates as the difference between the testing dimensionality and training dimensionality increases, as empirically revealed by [30]. For instance, PtrNet trained on 150 dimensions, when used to solve a problem of 1000 dimensions, has a disparity of up to 41.44% compared to the solution obtained by LKH [15]. As suggested in these works, to solve 1000-dimensional TSPs, it is more natural and effective to train PtrNet on the same dimensionality. Nevertheless, Table 1 illustrates that current methods predominantly focus on various COPs around 150 dimensions or less. While a few methods attempt solving 1000-dimensional COPs, their training instances are still no more than 150 dimensions, and the results all faced the issue of low quality. This situation remains the same for recently proposed Attention and POMO models, which underscores the importance of training and solving at similar dimensions to enhance the solution quality for high-dimensional COPs. However, this will result in more local optima in the parameter space and lead to a higher risk that the search process is trapped [31]. On this basis, this challenge is a natural scenario for verifying the merits of EAs for training PtrNet.
@@
  • We test the computational time needed by all algorithms to reach the same targeted solution quality on 1000-dimensional TSP datasets. It is found that PtrNet-EA is able to consume much fewer computational time than the compared algorithms.
@@
  • This is the first work empirically assesses the advantages of training PtrNets with EA over traditional gradient based training methods. This provides new insights to improve neural solvers with EAs.
  • This is the first work empirically verifies that training PtrNet on 1000-dimensional instances can significantly benefit the inference on 1000-dimensional instances. This suggests that scaling up training instances is in need to improve the neural solvers on large-scale problems.
@@ 3.2 Framework of PtrNet-EA
An evolutionary framework for NNs is an optimization approach that draws inspiration from the natural process of evolution, aiming to optimize the structure and/or parameters of NNs through iterative generations. For neural solvers addressing COPs, their structure remains constant during training, with only the network weights requiring adjustment. Based on this, in this work the evolutionary framework is employed to optimize the parameters of the PtrNet while keeping the architecture fixed. The following workflow describes how EAs optimize the parameters, each individual in the population denotes the parameters of a PtrNet.
@@ 3.3 Concrete EA-based Training Algorithm
In this paper, we choose NCS to instantiate the PtrNet-EA. NCS is a type of EA with explicitly enhanced diversity and has been applied to solve numerous practical problems [28, 40]. NCS consists of multiple parallel search processes. Each search process is driven by a Gaussian mutation operator [44], which generates the new individual $x_{i}^{\prime}$ by sampling from $\mathcal{N}\left(x_{i},\sigma_{i}\right)$, where $\sigma_{i}$ is the standard deviation of the Gaussian distribution $\mathcal{N}$ and represents the search step-size. The individuals are expected to be with high quality and distant from each other to form a parallel exploration of multiple regions of the parameter space. In NCS, the distance is defined as Bhattacharyya distance [45] by Eq.(1), where $p_{i}$ represents the distributions corresponding to other individuals. NCS forces the population seeking different optimal areas by trading off between the distance and the objective fitness values with an adaptive parameter $\lambda_{t}$. The $\sigma_{i}$ is updated based on the 1/5 success rule [46]. That is, $\sigma_{i}$ will be increased to enhance the exploration if 20% of the offspring $x_{i}^{\prime}$ is better than its parent $x_{i}$ within $t_{e}$ iteration; otherwise, it will be reduced for favoring exploitation, according to Eq.(2), where $c_{i}$ counts the replaced individuals. The pseudo-code is given in Algorithm , and at the end of the algorithm, the final population obtains all the PtrNet individuals with good performance and diversity. For more details of NCS, please refer to [27].
@@ 4.2 Algorithms Settings
Suggested by [47], three types of well-established combinatorial solvers are selected as the compared algorithms. The first type contains the PtrNet-based methods, including the original PtrNet trained by SL (PtrNet-SL), the Graph Pointer Network with 2-opt (GPN-RL) [15], the pointer network trained by RL (PtrNet-RL) [18]. The second type includes two non-PtrNet learning-based solvers, i.e., the Structure to Vector Deep Q-learning (S2V-DQN) [48] and Variable Strategy Reinforced LKH (VSR-LKH) [49]. S2V-DQN is a deep learning architecture over graphs, and VSR-LKH is a hybrid of RL and LKH. The third type consists of the heuristic search based SOTA methods, i.e., the 2-opt [50], the Google OR-Tools with saving algorithm, and Focused Ant Colony Optimization (FACO)[51].
@@ 4.3 Experimental Settings
This work aims to study how to train PtrNets with 1000-dimensionality TSP efficiently. We set the time budget to 500 minutes for all the algorithms for fairness. Indeed, any other time budgets larger than 500 minutes do not mean unacceptable. We simply approximate it around one working day of 8 hours for the consideration of real-world activities. More specifically, for learning-based solvers, the training phase is set to 500 minutes, as the testing phase is usually very efficient and thus the testing time can be omitted. For the heuristic search-based methods, we allow them to solve the instances in 500 minutes. Unless specified, all experimental results are the average of three independent runs. All experiments are conducted on a workstation with 320GB RAM and 96 cores (4.0GHz, 71.5MB L3 Cache), running Ubuntu 20.04. The experiments are divided into four Groups and will be described in detail.
Group I. We compare the proposed PtrNet-EA with the 8 chosen algorithms. All the learning-based solvers are first trained with 1 million training instances of TSP100, and then tested on 2000 testing instances of TSP100, TSP500, and TSP1000, respectively. The 2-opt, OR-Tools and FACO are directly used to solve the 2000 testing instances, as they do not involve the training step. The output path lengths of 2000 testing instances are averaged as the final performance of each algorithm.
@@ 4.4 Results and Analysis
Group I. When compared with PtrNet-SL, PtrNet-EA has 18.35%, 18.70% and 21.61% improvements on TSP100, TSP500 and TSP1000, respectively. When compared with RL-based PtrNets, PtrNet-EA has at least 11.95%, 17.23% and 10.91% improvements, respectively. These comparisons show that the PtrNet-EA has a great learning ability. When compared with other SOTA heuristic methods, the results have been improved by 4.40%, 4.27% and 2.22%, , which means that PtrNet-EA is able to achieve better solution in time-constrained tasks. The detailed results are shown in Table 4.
@@
Group II. The results are shown in Fig.4. It is obvious that the compared methods need over 10x the time budget to achieve the same performance of PtrNet-EA, and fully illustrates the efficiency of EA in training PtrNets.
Figure 4: 10x time budget for compared methods training on TSP1000 to reach the PtrNet-EA’s performance in 1x time budget.
@@
Table 6: Performance difference between PtrNet-EA and PtrNet-EA-P.
Group IV. Last but not least, the policy portfolio can also make some contributions. When we use all policies in the last population as a portfolio, each policy has the potential to obtain the optimal solution on specific problem instances. As shown in Fig.6, 34% best solutions of the 2000 testing instances are obtained by the policies in the portfolio other than the best policy. As much as 1.99% improvement is introduced by the policy portfolio (see Table 6), and the overall improvement reached an astonishing 30.21% compared with the original PtrNet-SL.
@@ 5 Conclusion
Therefore, our work provides a new insight to both fields of EAs and PtrNets, which indicating that EAs could train PtrNets more proficiently than gradient-based methods, e.g., SL and RL. This suggests that for applications that leverage PtrNet in solving large-scale COPs, utilizing EAs for training should be seriously considered.
Moreover, while our current research focused on the PtrNet, given the sequential structures also shared by the models like Attention and POMO, it is reasonable to induce that they can also benefit from similar evolutionary training framework. In the future, how to train models better with EAs will be studied as they appear more complex.

### 11 Mar 2024 v4

@@ Abstract
Pointer Network (PtrNet) is a specific neural network for solving Combinatorial Optimization Problems (COPs). While PtrNets offer real-time feed-forward inference for complex COPs instances, its quality of the results tends to be less satisfactory. One possible reason is that such issue suffers from the lack of global search ability of the gradient descent, which is frequently employed in traditional PtrNet training methods including both supervised learning and reinforcement learning. To improve the performance of PtrNet, this paper delves deeply into the advantages of training PtrNet with Evolutionary Algorithms (EAs), which have been widely acknowledged for not easily getting trapped by local optima. Extensive empirical studies based on the Travelling Salesman Problem (TSP) have been conducted. Results demonstrate that PtrNet trained with EA can consistently perform much better inference results than eight state-of-the-art methods on various problem scales. Compared with gradient descent based PtrNet training methods, EA achieves up to 30.21% improvement in quality of the solution with the same computational time. With this advantage, this paper is able to report the results of solving 1000-dimensional TSPs by training a PtrNet on the same dimensionality. This strongly suggests that the pointer network can be trained better with evolutionary algorithms for solving COPs, especially on high-dimensional instances.
@@ 1 Introduction
On the other hand, one major drawback of PtrNet is the lower quality of the results compared with traditional methods, especially on larger-scale COPs, as systematically evaluated by [13]. For example, the quality gap of result between PtrNet and the well-established solver Lin-Kernighan heuristic (LKH) [14] degenerate significantly from 19.81% to 41.44% on TSP with 250 nodes and 1000 nodes [15], respectively. Although recent works have proposed more advanced models, like Policy Optimization with Multiple Optima (POMO)[16], still cannot resolve the above issue. On this basis, this paper focuses on the training method of the neural networks (NNs) for COPs, rather than the model design. We initially opted for the PtrNet over the more complex models for simplicity and generality. The straightforward architecture of PtrNet provides a clear foundation for implementing and understanding the effects of EAs in training NNs, and establishes a baseline before exploring more intricate models.
@@
In order to verify the aforementioned advantages of EAs, we have designed sophisticated computational experiments on training PtrNet for TSPs. The experiments specially focus on an important training challenge. That is, existing training methods are hardly to train PtrNet effectively on 1000-dimensional instances, which are more realistic to real-world problem scales[28]. Albeit PtrNet is able to handle varied dimensional instances on its recurrent network structure, the quality of solutions deteriorates as the difference between the testing dimensionality and training dimensionality increases, as empirically revealed by [29]. We also found similar phenomenon by comparing the increasing performance gap between PtrNet and LKH on solving TSPs with 250, 500, 750, 1000 dimensions, where the PtrNet is trained by gradient descent on 50 dimensional instances. An empirical experiment shows PtrNet trained on 150 dimensions, when used to solve a problem of 1000 dimensions, has a disparity of up to 41.44% compared to the solution obtained by LKH [15]. As suggested in these works, to solve 1000-dimensional TSPs, it is more natural and effective to train PtrNet on the same dimensionality.
Nevertheless, Table 1 illustrates that current methods predominantly focus on various COPs around 150 dimensions or less. While a few methods attempt solving 1000-dimensional COPs, their training instances are still no more than 150 dimensions, and the results all faced the issue of low quality. This situation remains the same for recently proposed POMO models, which underscores the importance of training and solving at similar dimensions to enhance the solution quality for high-dimensional COPs. However, this will result in more local optima in the parameter space and lead to a higher risk that the search process is trapped[30]. On this basis, the challenge is a natural scenario for verifying the merits of EAs for training PtrNet.
@@
  • We test the computational time needed by all algorithms to reach the same targeted solution quality on 1000-dimensional TSP datasets. It is found that PtrNet-EA is able to consume much fewer computational time than the compared algorithms. All compared algorithms are accelerated with parallel computing units, if possible.
@@
  • This is the first work empirically assesses the advantages of training PtrNets with EA over traditional gradient based training methods by achieving in total 30.21% improvements on 1000-dimensional COPs. This provides new insights to train neural combinatorial solvers.
  • The proposed training framework help PtrNet outperform OR-Tools and 2-opt significantly, mitigating the criticisms that neural combinatorial solvers are not competitive to existing mathematic programming based or heuristic-based solvers.
  • It is verified via extensive experiments that the advantages of training PtrNets with EA lies in the diversified search ability and the parallel computing capability, clarifying the key points of amplifying the performance of EA as the PtrNet trainer.
The reminder of this paper is organized as follows. In section II, the related works are reviewed. The PtrNet-EA framework is presented in section III as well as its concrete version by employing the NCS algorithm. Sophisticated empirical studies are conducted in section IV to support the claims that EA is able to train PtrNet better. Conclusions are drawn in section V.
@@ 3.2 Framework of PtrNet-EA
An evolutionary framework for NNs is an optimization approach that draws inspiration from the natural process of evolution, aiming to optimize the structure and/or parameters of NNs through iterative generations [43]. For neural solvers addressing COPs, their structure remains constant during training, with only the network weights requiring adjustment. Based on this, in this work the evolutionary framework is employed to optimize the parameters of the PtrNet while keeping the architecture fixed. The following workflow describes how EAs optimize the parameters, each individual in the population denotes the parameters of a PtrNet.
@@ 3.3 Concrete EA-based Training Algorithm
In this paper, we choose NCS to instantiate the PtrNet-EA. NCS is a type of EA with explicitly enhanced diversity and has been applied to solve numerous practical problems [27, 39]. NCS consists of multiple parallel search processes. Each search process is driven by a Gaussian mutation operator [44], which generates the new individual $x_{i}^{\prime}$ by sampling from $\mathcal{N}\left(x_{i},\sigma_{i}\right)$, where $\sigma_{i}$ is the standard deviation of the Gaussian distribution $\mathcal{N}$ and represents the search step-size. The individuals are expected to be with high quality and distant from each other to form a parallel exploration of multiple regions of the parameter space. In NCS, the distance is defined as Bhattacharyya distance [45] by Eq.(1), where $p_{i}$ represents the distributions corresponding to other individuals. NCS forces the population seeking different optimal areas by trading off between the distance and the objective fitness values with an adaptive parameter $\lambda_{t}$. The $\sigma_{i}$ is updated based on the 1/5 success rule [46]. That is, $\sigma_{i}$ will be increased to enhance the exploration if 20% of the offspring $x_{i}^{\prime}$ is better than its parent $x_{i}$ within $t_{e}$ iteration; otherwise, it will be reduced for favoring exploitation, according to Eq.(2), where $c_{i}$ counts the replaced individuals. The pseudo-code is given in Algorithm 1 , and at the end of the algorithm, the final population obtains all the PtrNet individuals with good performance and diversity. For more details of NCS, please refer to [26].
@@
$\sigma_{i}=\begin{cases}\sigma_{i}/0.99,\;\text{if }c_{i}/t_{e}\leq 0.2;\\ 0.99\sigma_{i},\;\text{otherwise}.\end{cases}$ (2)
Algorithm 1 PtrNet-EA
Input: $T_{max}$,$t_{e}$,$N$
Output: Last population
1:  Randomly generate an initial population of N PtrNets.
2:  Evaluate the PtrNets by objective function $f$.
3:  Set $t\leftarrow 0$.
4:  while $t<T_{max}$ do
5:     Set $\lambda_{t}\leftarrow\mathcal{N}(1,0.1-0.1t/T_{max})$
6:     for  $i=1$ to $N$ do
7:        Generate a new PtrNet $x_{i}^{\prime}\leftarrow\mathcal{N}(x_{i},\sigma_{i})$
8:        Compute $f(x^{\prime}_{i})$ and $Corr(p_{i}^{\prime})$
9:        $x_{i}\leftarrow x_{i}^{\prime}$ if $f(x^{\prime}_{i})$ / $Corr(p_{i}^{\prime})<\lambda_{t}$
10:     end for
11:     $t\leftarrow t+1$
12:     if $mod(t,t_{e})=0$ then
13:        Update every $\sigma_{i}$ according to Eq.(2)
14:     end if
15:  end while
16:  return solution
@@ 4.2 Algorithms Settings
Suggested by [47], three types of well-established combinatorial solvers are selected as the compared algorithms. The first type contains the PtrNet-based methods, including the original PtrNet trained by SL (PtrNet-SL), the Graph Pointer Network with 2-opt (GPN-RL) [15], the pointer network trained by RL (PtrNet-RL) [17]. The second type includes three non-PtrNet learning-based solvers, i.e., POMO[16], the Structure to Vector Deep Q-learning (S2V-DQN) [48] and Variable Strategy Reinforced LKH (VSR-LKH) [49]. S2V-DQN is a deep learning architecture over graphs, and VSR-LKH is a hybrid of RL and LKH. The third type consists of the heuristic search based SOTA methods, i.e., the 2-opt [50], the Google OR-Tools with saving algorithm, and Focused Ant Colony Optimization (FACO)[51].
@@ 4.3 Experimental Settings
This work aims to study how to train PtrNets with 1000-dimensionality TSP efficiently. We set the time budget to 500 minutes for all the algorithms for fairness. Indeed, any other time budgets larger than 500 minutes do not mean unacceptable. We simply approximate it around one working day of 8 hours for the consideration of real-world activities. More specifically, for learning-based solvers, the training phase is set to 500 minutes, as the testing phase is usually very efficient and thus the testing time can be omitted. For the heuristic search-based methods, we allow them to solve the instances in 500 minutes. Unless specified, all experimental results are the average of three independent runs on the target TSP testing set each with 2000 problem instances. All experiments are conducted on a workstation with 320GB RAM and 96 cores (4.0GHz, 71.5MB L3 Cache), running Ubuntu 20.04111The code can be found https://github.com/mythezone/PtrNetEA.git. The experiments are divided into four Groups and will be described in detail.
Group I. We compare the proposed PtrNet-EA with the 9 chosen algorithms. All the learning-based solvers are first trained with 1 million training instances of TSP100, and then tested on 2000 testing instances of TSP100, TSP500, and TSP1000, respectively. The 2-opt, OR-Tools and FACO are directly used to solve the 2000 testing instances, as they do not involve the training step. The output path lengths of 2000 testing instances are averaged as the final performance of each algorithm.
@@ 4.4 Results and Analysis
Figure 4: A comparison of the convergence curves during the training process of other methods against PtrNet-EA. The three columns correspond to the training conducted on TSP100, TSP500, and TSP1000, respectively. And the three rows depict three groups of different compared algorithms. The vertical axis represents the Tour Length for the current batch on the Training set.
Group I. As illustrated in Fig.4, throughout the training process, the PtrNet-EA not only demonstrates a faster convergence rate but also achieves the best performance on the training set among all the evaluated methods.
Comparing the training curve across three columns in Fig.4, it is observed that in TSP100 and TSP500, PtrNet-EA does not show a significant difference in convergence trends compared to other methods. However, as the problem size increases to TSP1000, the disparity between PtrNet-EA and other approaches becomes more pronounced. The first row of comparative reveals that, across TSP100, TSP500, TSP1000, the EA trains PtrNet more effectively, when compared with SL and RL. Notably, since the PtrNet-SL method employs approximate solutions as labels for training, it converges faster than the RL method in some cases. Comparing the second and third rows, on the training set, the EA achieves better training performance. This highlights the effectiveness of PtrNet-EA in terms of both learning efficiency and solution quality within the given training scope.
When compared with PtrNet-SL, PtrNet-EA has 18.35%, 18.70% and 21.61% improvements on TSP100, TSP500 and TSP1000, respectively. When compared with PtrNet-RL, PtrNet-EA has at least 11.95%, 17.23% and 10.91% improvements on TSP100, TSP500 and TSP1000, respectively. These comparisons show that the PtrNet-EA has a great learning ability. When compared with other SOTA methods, the results have been improved by at least 4.40%, 4.27% and 2.22% (FACO) on TSP100, TSP500, TSP1000, respectively. The result means that PtrNet-EA is able to achieve better solution in time-constrained tasks. The detailed results are shown in Table 4.
@@
Integrating the results of these two comparative experiments in Group I, it can be inferred that with the premise of a reasonable time budget, EAs demonstrate superior efficacy in searching the parameter space to identify optimal solutions, compared to other methods. Furthermore, when compared to traditional methods, PtrNet-EA achieves better results within the same computational time budget. Therefore, we could draw the conclusion that under the constraint of limited computational resources, the EA method trains neural networks more effectively and yields superior solutions.
@@
Group II. The results are shown in Fig.5. It is obvious that the compared methods need over 6.5x the time budget to achieve the same performance of PtrNet-EA. PtrNet-SL and PtrNet-RL even cannot achieve the same performance of PtrNet-EA in 10x the time budget.
We posit that the observed results are attributable to two primary factors. Firstly, the population-based search characteristic of EAs aligns well with the highly parallelized architecture of modern computers, enabling more efficient utilization of hardware for acceleration compared to other algorithms. Secondly, the selection of an EA that is well-matched to the problem (in our case, training neural networks) results in a more effective task completion.
NCS explicitly incorporates the distance between individuals in the population into the fitness function and adaptively adjusts the step size during the searching process. This approach emphasizes population diversity and the characteristic of obtaining offspring through sampling distributions, which makes NCS particularly suitable for searching high-dimensional neural network parameters. It also suggests that EAs with good diversity maintenance techniques should be able to train PtrNet well.
Figure 5: 10x time budget for compared methods training on TSP1000 to reach the PtrNet-EA’s performance training in 500 minutes.
@@
This set of experiments comprehensively demonstrates the superiority of EAs in training neural networks. Under the condition of maintaining diversity among individuals in the population, a larger population size signifies a more effective exploration of the parameter space. This not only implies the potential for a time-efficient search by trading resources for time, achieving the same solution quality in less time. It also suggests the possibility of obtaining better results within the same time budget by further increasing computational resources. In the context of distributed computation architectures, where abundant and cost-effective computing resources are readily available, a method that can fully leverage these resources will offer more superior solutions for users with specific needs.
@@
Table 6: Performance difference between PtrNet-EA and PtrNet-EA-P.
@@
Group IV. Last but not least, the policy portfolio can also make contributions to the testing performance. When we use all policies in the last population as a portfolio, each policy has the potential to obtain the optimal solution on specific problem instances. As shown in Fig.7, 34% best solutions of the 2000 testing instances are obtained by the policies in the portfolio other than the best policy. As much as 1.99% improvement is introduced by the policy portfolio (see Table 6), and the overall improvement reached an astonishing 30.21% compared with the original PtrNet-SL.
While other stochastic methods may also employ portfolio polices to solve problems multiple times, retain the best solution to enhance performance without adding extra waiting time due to parallel execution. The solutions obtained might lack diversity, which means the results of each run could be very similar. In contrast, the EAs inherently comprises population with multiple individuals. It not only avoids additional time budget but also utilizes the diversity within these individuals with minimal computational resources required for infering the solutions. This approach enables the generation of substantially varied computational outcomes, thereby potentially improving the quality of the solutions.
@@ 5 Conclusion
Therefore, our work provides a new insight to both fields of EAs and neural combinatorial solvers, which indicating that EAs could train PtrNets more proficiently than gradient-based methods. This suggests that for applications that leverage PtrNet in solving large-scale COPs, utilizing EAs for training should be seriously considered.
Moreover, while our current research focused on the PtrNet, it is reasonable to induce that the other neural combinatorial solvers can also benefit from similar evolutionary training framework. In the future, how to train more advanced solvers with EAs will be studied as their network models appear to be more complex.